OpenStudy (anonymous):

Answer as a=fasdt as yuou can!!! Mary has $4.00 in nickels, dimes, and quarters. If she has twice as many dimes as quarters and forty-seven more nickels than dimes, how many coins of each type does she have?

OpenStudy (anonymous):

I'm sorry to be asking this, but I don't live in the u.s. so... how much of a dollar is a nickel and a dime? I suppose a quarter is 1/4 of a dollar.

OpenStudy (anonymous):

.05 nick, .10 dime

OpenStudy (anonymous):

Thank you. Let's see then: She has four total dollars. I'll call nickel 'n', dime 'd' and quarter 'q'. Twice as many dimes as quarters = > d = 2*q 47 more nickles than dimes = > n = d + 47 0.05 n + 0.10 d + 0.25 q = 4 0.05 (d + 47) + 0.10 d + 0.25 (d/2) = 4 0.05 d + 2.35 + 0.10 d + 0.125 d = 4 0.275 d = 4 - 2.35 0.275 d = 1.65 275 d = 1650 55 d = 330 11 d = 66 d = 6